We are heading toward the summer solstice and we had another great week in the field. The weather has been slightly cooler and more overcast this week in McMurdo and the dry valleys. We had a few weather-related delays and aborted missions with helicopter operations. This shift in weather resulted in less stream flow, but slightly improving conditions for working on the lakes.
• C-504 (glaciers) Anna and Forrest completed the ice core sampling on Canada Glacier and will continue with sampling the glacier melt.
• C-505 (limno) is well into L2. They completed the second limno runs at Lake Fryxell and Lake Miers. They are conducting the limno run at Lake Hoare today and plan to move to Lake Bonney on Tuesday.
• C-506 (streams) The stream team continues their summer routine of site visits and water sampling.
• C-509 (geochem) has been analyzing water samples on the IC as they are received from the field. Kathy has also been setting up the new Shimadzu instrument in lab 107 with help from Nik.
• C-511 (met + lake physical) Maciek reports that they completed a successful field season. The team returned to McMurdo yesterday (Saturday). All Blue Boxes and met stations were serviced and are working. Blue Boxes in Taylor Valley were upgraded to a hashtag design; Miers will be upgraded next season. Telemetry in the western end of TV was upgraded with a repeater installation on Bonney Riegel (installation done by Thomas) allowing both LB BBs to have more reliable comms. Lastly, C-511 build an independent UV monitor station (for a year-round data collection) for Anna/Mike project and installed it on the LH met station tripod before leaving the field.
The Good -
• Overall, things are going well. We continue to get great support from ASC, especially Rae & Renee at Lake Hoare, Lindsay for scheduling helos, and the rest of the helo crew.
Nothing to report.
Nothing to report.